OpenStudy (anonymous):

You select a card at random. Without replacing the card, you select a second card. Find the probability. M A T H E M A T I C S P(M, then H) A. 3 /11 B. 2/21 C. 1/55 D. 2/121

Miracrown (miracrown):

is there more information to this problem? for example what kind of cards are being selected?

Miracrown (miracrown):

what is on the cards

OpenStudy (anonymous):

M A T H E M A T I C S

Miracrown (miracrown):

ok so there are a total of 11 cards? each one has a letter that makes up the word MATHEMATICS?

OpenStudy (anonymous):

Yes

Miracrown (miracrown):

ok so we need to find the probability of selecting an M and then an H

Miracrown (miracrown):

ok so what is the probability of selecting an M on the first draw?

OpenStudy (anonymous):

I think that the question. I can't pin point it

Miracrown (miracrown):

it will be equal to the number of cards with an M divided by the total number of cards

Miracrown (miracrown):

ok so how many cards have an M? M A T H E M A T I C S there is one card for each letter

OpenStudy (anonymous):

2

OpenStudy (anonymous):

P(draw an M) = (number favorable to drawing an M) / (number of cards total), to re-phrrase Miracrown = \frac{2}{11}

OpenStudy (anonymous):

\[\frac{2}{11}\]

Miracrown (miracrown):

yyes and there are a total of 11 cards so the probability of choosing an M as the first card is 2/11 now we need to find the probability of choosing an H when we choose a second card

Miracrown (miracrown):

How many H cards are there? 1 right? right now remember we don't replace the cards once we have chosen them, so there are 10 cards left since we chose one already for the first draw so the probability of choosing an H on the second draw is 1/10 since there are 10 cards left, only 1 is an H so the probability of selecting an M as the first card and an H as the second is 2/11 times 1/10

Miracrown (miracrown):

so the probability of selecting an M as the first card and an H as the second is 2/11 times 1/10 remember P(A and B) = P(A) times P(B) when the events A and B are independent as they are here here A = choosing an M as the first card B = choosing H as the second given we chose an M as the first

Miracrown (miracrown):

actually they arent really independent, but we can still use that formula

Miracrown (miracrown):

2/11 times 1/10 = 11/55 1/55

OpenStudy (anonymous):

Alright! Thank you >_<

Miracrown (miracrown):

so the answer is C

Miracrown (miracrown):

yw :]

OpenStudy (anonymous):

\[P(AB) = P(B)P(B|A)\] works for independent and dependent (if independent P(B given A) = P(B)

OpenStudy (anonymous):

^^ only formula to memorize cuz it applies to both

OpenStudy (anonymous):

sorry that should be \[P(A)P(B|A)\]